OpenStudy (anonymous):

The half-life of Radium-226 is 1590 years. If a sample contains 400 mg, how many mg will remain after 3000 years?

OpenStudy (anonymous):

\[\large 400\times \left(\frac{1}{2}\right)^{\frac{3000}{1590}}\] and a calculator